help solving this equation
2/5(15x+35)-3=-1/2(12x-6)

Answer by checkley77(12844)

2/5(15x+35)-3=-1/2(12x-6)
6x+14-3=-6x+3
6x+6x=3-14+3
12x=-8
x=-8/12
x=-2/3 ans.
Proof:
2/5(15*-2/3+35)-3=-1/2(12*-2/3-6)
2/5(-10+35)-3=-1/2(-8-6)
2/5(25)-3=-1/2*-14
10-3=7
7=7